OpenStudy (anonymous):

(d^2y)/(dx^2) of sqrt(y) + xy = 1. I can get the 1st implicit derivative, but I lose it somewhere on the 2nd.

OpenStudy (amistre64):

show the first one then

OpenStudy (anonymous):

for dy/dx I get - (y)/(1/2sqrt(y) + x)

OpenStudy (amistre64):

sqrt(y) + xy = 1 1/2sqrt(y) y' + x'y + xy' = 0 since x' = dx/dx in this case, that goes to 1 1/2sqrt(y) y' + y + xy' = 0 y'(1/2sqrt(y) + x) = -y y' = -y/(1/2sqrt(y) + x)

OpenStudy (shayaan_mustafa):

Hi Ajk :) How are you friend ?

OpenStudy (amistre64):

you can either go ahead from the unsimplified y'; or this one :)

OpenStudy (shayaan_mustafa):

@mahmit2012 I think he is asking for implicit differentiation. amistre64 is doing right.

OpenStudy (anonymous):